Selima deposits $8000 ( t = 0 years) in an account paying r(2) = 5.00%.At the end of 4 years (t = 4 years) she adds another $5000 to her account and the interest rate changes to r(2) = X%. After another 3 years ( t = 7 years) she adds another $7000 to her account and the interest changes to r(2) = 3.00%. At the end of 11 years, she has $27721.81 in her account.

How much money does she have in her account at t = 4 years just before she makes her deposit?

b) How much money does she have in her account after 7 years, just after she makes her deposit?

c) What is X%?
1
Expert's answer
2020-06-15T17:48:19-0400

a.Amount befor the second deposit is


"P(1+r)^n"


"8000(1+{5\over100})^4"


"=9724.05"

b.

Amount after 3yrs at t=7


"9724.05+50000=14724.05=p"


"14724.05(1+x)^7"


c.

Value of x


"14724.05(1+x)^7+7000(1+{3\over100})^{11}=27721.81"




"14724.05(1+x)^7+9689.63709=27721.81"


"14724.05(1+x)^7=18032.0.179"


"(1+x)^7=1.2474"


"1.0321=1+x"


"x=0.0321=3.21\%"